# Patrick Cheridito and Moritz Weiss

### v1

@@ Appendix D Feature normalization
  • We recall from Section 3.1 that the algorithm observes the first $K-1$ entries of the order book volumes defined in equation (1). Instead of the absolute, we consider the normalized volumes on the bid and ask sides $\bar{v}^{b}(t)\in\operatorname{\mathbb{R}}^{K-1}_{+}$ and $\bar{v}^{a}(t)\in\operatorname{\mathbb{R}}_{+}^{K-1}.$ We normalize the volumes by using the long-term average queue sizes $\tilde{v},$ defined in Appendix A.2, as follows:
@@
  • We normalize the market order flow $\Delta m(t)$ in $(t-\Delta t,t]$ by dividing it by the total volume of all market buy and sell orders in $(t-\Delta t,t]$. The normalized market order flow is then contained in $[-1,1].$
  • We normalize the limit order flow $\Delta l(t)$ in $(t-\Delta t,t]$ by dividing it by the total volume of all limit buy and sell orders in $(t-\Delta t,t]$. The normalized limit order flow is then contained in $[-1,1].$
@@
  • We normalize the inventory relative to the initial inventory by considering $M(t)/M\in[0,1].$
@@
  • If there are $m\leq M(t)$ active limit orders in the book, and $u\leq M(t)$ is the number of lots that are withheld from the order book, and the remaining orders have already been filled, then we encode the levels of all limit orders by
@@
  • If there are $m\leq M(t)$ active limit orders in the book, and $u\leq M(t)$ is the number of lots that are outside the book, and the remaining orders have already been filled, then we encode the queue positions of all limit orders by
@@
    Here, for $k\in\{1,\dots m\},$ the quantity $q^{k}$ is the queue position of the active limit orders, ordered from the smallest to the highest price levels and queue positions. For $k\in\{m+1,\dots,m+u\},$ we set $q^{k}=50.$ For $k\in\{m+u+1,\dots,M\},$ we set $q^{k}=-50.$ In our simulations, the queue sizes are rarely larger than $50$ (see Figures 5(a) and 5(b)), which justifies the normalization of the queue positions by $50$
@@
  • We use one additional feature, representing the number of active orders per price level, given by $\gamma(t)\in[0,1]^{K-1},$ where $K$ is the dimension of the simplex containing the actions. For $k\in\{1,\dots,K-1\},$ the quantity $\gamma^{k}(t)$ is the fraction of the inventory $M(t)$ that is placed at $k-1$ ticks above the best ask and $\gamma^{K}(t)$ is the fraction of the inventory $M(t)$ that is either placed higher than $K-2$ ticks away from the best ask price or not placed in the book. We note that the feature $\gamma(t)$ is redundant since the algorithm already has information on the levels of its limit orders through the vector $\bar{l}(t).$ However, we noticed that the inclusion of $\gamma(t)$ improves the convergence speed of the algorithm.

### v2

@@ B.3 Feature normalization
  • We recall from Section 3.1 that the algorithm observes the first K1K-1 entries of the order book volumes defined in equation (1). Instead of the absolute, we consider the normalized volumes on the bid and ask sides $\bar{v}^{b}(t)\in\operatorname{\mathbb{R}}^{K-1}_{+}$ and $\bar{v}^{a}(t)\in\operatorname{\mathbb{R}}_{+}^{K-1}.$ We normalize the volumes by using the long-term average queue sizes $\tilde{v}$, defined in Appendix A.2, as follows:
@@
  • We normalize the market order flow $\Delta^{M}(t)$ in $(t-\Delta t,t]$ by dividing it by the total volume of market buy and sell orders in $(t-\Delta t,t]$. Then, the normalized market order flow is contained in [1,1].[-1,1].
  • We normalize the limit order flow $\Delta^{L}(t)$ in $(t-\Delta t,t]$ by dividing it by the total volume of limit buy and sell orders in $(t-\Delta t,t]$. Then, the normalized limit order flow is contained in [1,1].[-1,1].
  • We normalize the cancellation order flow $\Delta^{C}(t)$ in $(t-\Delta t,t]$ by dividing it by the total volume of limit sell and buy order cancellations in $(t-\Delta t,t]$. Then, the normalized cancellation order flow is contained in [1,1].[-1,1].
@@
  • We normalize the inventory relative to the initial position by considering $M(t)/M\in[0,1].$
@@
  • If there are $m\leq M(t)$ active limit orders in the book, and $u\leq M(t)$ is the number of lots that are inactive, and the remaining orders have already been filled, then we encode the levels of all limit orders by
@@
  • If there are $m\leq M(t)$ active limit orders in the book, and $u\leq M(t)$ is the number of lots that are inactive, and the remaining orders have already been filled, then we encode the queue positions of all limit orders by
@@
    Here, for $k\in\{1,\dots m\},$ the quantity $q^{k}$ is the queue position of the active limit orders, ordered from the smallest to the highest price levels and queue positions. For $k\in\{m+1,\dots,m+u\},$ we set $q^{k}=50.$ For $k\in\{m+u+1,\dots,M\},$ we set $q^{k}=-50.$ In our simulations, the queue sizes are rarely larger than $50$ (see Figures 8(a) and 8(b)), which justifies the normalization of the queue positions by $50.$
Additional feature
@@
We use one additional feature, representing the number of active orders per price level, given by $\gamma(t)\in[0,1]^{K-1},$ where $K$ is the dimension of the simplex containing the actions. For $k\in\{1,\dots,K-1\},$ the quantity $\gamma^{k}(t)$ is the fraction of the inventory $M(t)$ that is placed at k1k-1 ticks above the best ask and $\gamma^{K}(t)$ is the fraction of the inventory $M(t)$ that is either placed higher than K1K-1 ticks away from the best ask price or not placed in the book. We note that the feature $\gamma(t)$ is redundant since the algorithm already has information on the levels of its limit orders through the vector $\bar{l}(t).$ However, the inclusion of $\gamma(t)$ improves the convergence speed of the algorithm. Furthermore, we need to compute $\gamma(t)$ to implement the new order allocation of the algorithm after choosing action $a(t)$, so there are no added computational resources required to compute this feature.
Appendix C Additional experiments
C.1 Generalized advantage function estimates
In equation (15), we use a version of the generalized advantage function estimate (GAE), which was introduced in Schulman et al. (2015). The GAE depends on two parameters: The discount factor $\gamma^{\text{GAE}}$ and a parameter $\lambda^{\text{GAE}}\in(0,1).$ The parameters are called $\gamma$ and $\lambda$ respectively in Schulman et al. (2015). For the trade execution problem, we have $\gamma^{\text{GAE}}=1.$ The parameter $\lambda^{\text{GAE}}$ controls the tradeoff between bias and variance of the advantage function estimate. A lower value of $\lambda^{\text{GAE}}$ leads to a low-variance/high-bias estimate. A high value for $\lambda^{\text{GAE}}$ leads to a high-variance/low-bias estimate. In (15), we choose a high-variance/low-bias estimate by setting $\lambda^{\text{GAE}}=1.$ We control the variance by choosing a larger number of trajectories $\tau=1,280$ per gradient step. In this section, we evaluate whether it is possible to achieve similar performance by choosing a smaller number of trajectories $\tau$ while controlling variance with a smaller value for $\lambda^{\text{GAE}}.$ In Schulman et al. (2015), the authors recommend a choice of $\lambda^{\text{GAE}}\in[0.9,0.99].$
We conduct two experiments with a smaller number of trajectories $\tau=320.$ We choose $\lambda^{\text{GAE}}=0.95$ to reduce variance and compare it with the baseline $\lambda^{\text{GAE}}=1.0$. The results are displayed in Table 8. In both cases, the performance of the algorithm is worse than in the case where it is using $\tau=1,280$ trajectories; see Table 1. Furthermore, we observe that setting $\lambda^{\text{GAE}}=0.95$ does not significantly improve the algorithm’s performance. An explanation for this might be that we are using only 10 time steps. In cases where the execution problem lasts over more time steps (like 100 or 1000 time steps), the variance of the advantage function estimate (15) will be larger, and choosing a smaller value for $\lambda^{\text{GAE}}=0.95$ might be beneficial.
Table 8: Results when the algorithm is trained with $\tau=320$ trajectories per gradient step, using $\lambda^{\text{GAE}}=1.0$ and $\lambda^{\text{GAE}}=0.95.$ The results for $\lambda^{\text{GAE}}=1.0$ are displayed in the upper panel of the table, while the results for $\lambda^{\text{GAE}}=0.95$ are displayed in the lower panel. The best expected values across both values for $\lambda^{\text{GAE}}$ are highlighted in bold. If two expected values are equal, we highlight both of them.
Market #Lots $\mathbb{E}[\text{SL}]$ $\sigma[\text{SL}]$ $\mathbb{E}[\text{TWAP}]$ $\sigma[\text{TWAP}]$ $\mathbb{E}[\text{LN}]$ $\sigma[\text{LN}]$
$\lambda^{\text{GAE}}=1.0$
Noise 20 0.52 1.19 -0.06 0.94 0.53 1.1
60 -1.09 1.34 -1.4 0.98 -0.78 0.89
Noise & Tactical 20 0.1 1.43 0.48 0.68 0.74 0.64
60 -3.36 0.99 -0.96 0.95 -0.36 0.64
Noise & Tactical 20 -1.64 2.95 -0.36 3.03 1.1 2.14
& Strategic 60 -2.51 3.67 -1.45 3.46 -0.01 2.39
$\lambda^{\text{GAE}}=0.95$
Noise 20 0.52 1.19 -0.06 0.94 0.6 0.82
60 -1.09 1.34 -1.4 0.98 -0.78 0.86
Noise & Tactical 20 0.1 1.43 0.48 0.68 0.72 0.72
60 -3.36 0.99 -0.96 0.95 -0.34 0.66
Noise & Tactical 20 -1.64 2.95 -0.36 3.03 1.12 2.13
& Strategic 60 -2.51 3.67 -1.45 3.46 -0.24 2.38
C.2 Variance learning
In this section, rather than using the schedule (13), we let the logistic-normal algorithm learn variances. As in Section 4.3, let $\Sigma=(\Sigma^{ij})_{i,j=0,1,\dots,K-1}$ be the covariance matrix of the logistic-normal distribution. We set $\Sigma^{ij}=0$ for $i\neq j$ and encode the logarithms of the variances
$\log(\Sigma^{kk})=\varsigma^{k}\in\operatorname{\mathbb{R}},\quad k=0,1,\dots,K-1.$ (23)
We then treat the vector $(\varsigma^{0},\dots,\varsigma^{K-1})\in\operatorname{\mathbb{R}}^{K}$ as a learnable but state independent parameter. We initialized the log-variances to $\varsigma^{k}=0$ for $k=0,1,\dots,K-1$ at the beginning of the training cycle. In each step, the parameters $\varsigma^{k},$ for $k=0,\dots,K-1,$ are then updated as part of the policy updating step using the loss function (16). We could also consider state-dependent variances, but the authors in Andrychowicz et al. (2021) find that this does not yield better results.
We conducted an additional test, using the same parameter settings as above for the market simulation and the logistic-normal algorithm. When evaluating the LN algorithm with scheduled variance scaling, we use a deterministic version by using $h(\mu_{\theta})$ as the actions, as in Section 6.3. When evaluating the LN algorithm with learned action variances, we use the stochastic policy for evaluation by sampling from the logistic-normal distribution with mean $\mu_{\theta}$ and variance $(\exp(\varsigma^{0}),\exp(\varsigma^{1}),\dots,\exp(\varsigma^{K-1}).$
We find that the performance of the algorithm with manual variance scaling (LN) and learnable variances (LNVR) is similar across all cases except when the algorithm sells 60 lots in the market consisting of noise and tactical traders. In this case, we see better convergence of the average returns per batch in Figure 10(b) and significantly better performance in Table 9 of the LN algorithm compared to the LNVR algorithm. Therefore, we prefer to manually reduce action variances at each gradient step, but in new applications both approaches should be tested.
Table 9: The table shows returns of the LN algorithm and all benchmark algorithms. Additionally, we include the LNVR algorithm, which also learns the action variances. The best expected returns per row are highlighted in bold.
Market #Lots $\mathbb{E}[\text{SL}]$ $\sigma[\text{SL}]$ $\mathbb{E}[\text{TWAP}]$ $\sigma[\text{TWAP}]$ $\operatorname{\mathbb{E}}[\text{LNVR}]$ $\sigma[\text{LNVR}]$ $\mathbb{E}[\text{LN}]$ $\sigma[\text{LN}]$
Noise 20 0.52 1.19 -0.06 0.94 0.64 0.93 0.61 1.03
60 -1.09 1.34 -1.4 0.98 -0.77 0.99 -0.72 0.9
Noise & Tactical 20 0.1 1.43 0.48 0.68 0.74 0.7 0.81 0.64
60 -3.36 0.99 -0.96 0.95 -0.41 0.69 -0.25 0.67
Noise & Tactical 20 -1.64 2.95 -0.36 3.03 1.15 2.14 1.13 2.08
& Strategic 60 -2.51 3.67 -1.45 3.46 0.14 2.04 0.23 2.15
(a) Noise Traders
(b) Noise & Tactical Traders
(c) Noise, Tactical & Strategic Traders
Figure 9: Reward histograms for the heuristic benchmark strategies and RL algorithms. The upper panel shows the results for an initial position of 20 lots, whereas the lower panel shows those for an initial position of 60 lots. The symbol LN refers to the algorithm with manual variance scaling, and LNVR refers to the algorithm that learns action variances.
(a) Noise Traders
(b) Noise & Tactical Traders
(c) Noise, Tactical, & Strategic Traders
Figure 10: Average episode reward per batch during training for different environments and initial position sizes. The x-axis shows the number of gradient steps. The top row shows the rewards for 20 lots. The bottom row shows the rewards for 60 lots. The symbol LN denotes the logistic-normal algorithm. LNVR denotes the logistic-normal algorithm with learnable variance parameters.
C.3 Long time horizon
In this section, we test a longer execution horizon of $T=300s.$ Since the algorithm has more time to sell its inventory, we adjust the initial position sizes to 40 lots and 120 lots. To maintain the same number of decision steps per execution, we set the algorithms’ decision interval to $\Delta t=30s.$ All other simulation parameters remain the same as before.
The LN algorithm outperforms all other algorithms, as shown in Table 10. Figure 11 shows the reward histograms for all algorithms. The reward convergence plots in Figure 12 also demonstrate a more stable and faster convergence of the LN algorithm compared to the DR algorithm. In the market with strategic traders (Figure 12(c)), we observe that the DR algorithm’s rewards become unstable toward the end of the training cycle. This may be mitigated by clipping, as in PPO Schulman et al. (2017), but we leave this to future research.
Table 10: Returns of all heuristic benchmark algorithms and the logistic-normal and Dirichlet algorithm, for an initial position of 40 and 120 lots, and an execution horizon of 300 seconds. The best expected values per row are highlighted in bold.
Market #Lots $\mathbb{E}[\text{SL}]$ $\sigma[\text{SL}]$ $\mathbb{E}[\text{TWAP}]$ $\sigma[\text{TWAP}]$ $\mathbb{E}[\text{DR}]$ $\sigma[\text{DR}]$ $\mathbb{E}[\text{LN}]$ $\sigma[\text{LN}]$
Noise 20 -0.35 1.43 -0.58 1.1 0.27 1.38 0.3 1.4
60 -2.75 1.29 -3.02 1.06 -1.72 1.22 -1.64 1.21
Noise & Tactical 20 -2.0 1.43 -0.19 1.0 0.55 0.75 0.61 0.75
60 -5.57 0.77 -4.13 1.03 -0.74 0.78 -0.72 0.81
Noise & Tactical 20 -2.11 3.29 -1.09 3.28 1.62 3.61 2.32 3.69
& Strategic 60 -3.96 4.01 -3.42 4.18 -1.35 4.33 0.26 3.21
(a) Noise Traders
(b) Noise & Tactical Traders
(c) Noise, Tactical & Strategic Traders
Figure 11: Reward histograms for all heuristic benchmark strategies and the logistic-normal and Dirichlet algorithms, for an initial position of 40 and 120 lots, and an execution horizon of 300 seconds. The upper panels show the results for 40 lots. The lower panels show the results for 120 lots.
(a) Noise Traders
(b) Noise & Tactical Traders
(c) Noise, Tactical & Strategic Traders
Figure 12: Average episode returns per batch for an execution horizon of 300 seconds and initial positions of 40 and 120 lots. The x-axis shows the number of gradient steps, and the y-axis shows the average returns per batch. The upper panels correspond to an initial position of 40 lots, and the lower panel to 120 lots. LN denotes the logistic-normal algorithm, and DR the Dirichlet algorithm.
C.4 Feature importance
Finally, we conduct experiments to evaluate feature importance. We tested the LN algorithm across all markets and for all initial positions, while omitting certain features. More precisely, we train the algorithm in three cases. The first case uses all features (LN). The second case uses all features except those that reflect order price levels and queue positions, defined in the fourth item of the private states in Section 3.1, and is referred to as LN-O. The second case uses all features, except the order book volume features, defined in the second item of the market states in Section 3.1, and is denoted by LN-V. The third case uses all features, except the market, limit, and cancellation order flow imbalance, and the mid-price price drift feature, defined in the last four items of the market states in Section 3.1, and is denoted by LN-D.
The results are reported in Table 11 and Figure 13. The most important features are the order book volumes. Removing them leads to the worst performance across all markets and initial positions, except for an initial position of 60 lots in the market with noise and tactical traders, and for an initial position of 60 lots in the market with noise, tactical, and strategic traders. This is expected, as volume imbalance is a strong predictor of short-term price movements. Interestingly, the volume features remain important even in the market with strategic traders, where one might expect order-flow or mid-price drift information to dominate. This suggests that order book volumes also help detect the presence and direction of strategic traders. The order-flow and mid-price drift features are the second-most-important set of inputs, particularly in markets with strategic traders, where dropping them leads to substantially worse performance. They are less influential in the market with only noise traders, and in the market with noise and tactical traders, where using all features offers only marginal improvement. The least important features are those related to the levels and queue positions of the algorithm’s own orders. In the market with noise and tactical traders, when 60 lots are traded, omitting these features even yields slightly better performance, indicating that volume information dominates in this environment. However, in the market with strategic traders, removing these features significantly degrades performance. In this market, queue position information becomes important for execution timing. If the strategic trader is selling, the market is more likely to drift downward, making it advantageous for a seller to be near the front of the queue and secure a fill before prices fall. Conversely, if the strategic trader is buying and the market is likely to drift upward, being further back in the queue is preferable, as the order can still be canceled to capture a better selling price if the queue size at the best ask price decreases and an upward price move becomes more likely.
Table 11: Rewards for different RL algorithms, omitting certain features. LN corresponds to the logistic–normal algorithm that uses all features. LN-O denotes the algorithm without level and queue-position features. LN-V corresponds to the algorithm without volume features. LN-D corresponds to the algorithm without market, limit, and cancellation order imbalance features, and the mid-price drift feature. The best expected values per row are highlighted in bold.
Market #Lots $\mathbb{E}[\text{LN-D}]$ $\sigma[\text{LN-D}]$ $\mathbb{E}[\text{LN-V}]$ $\sigma[\text{LN-V}]$ $\mathbb{E}[\text{LN-O}]$ $\sigma[\text{LN-O}]$ $\mathbb{E}[\text{LN}]$ $\sigma[\text{LN}]$
Noise 20 0.61 1.04 0.38 0.94 0.61 1.01 0.61 1.03
60 -0.79 0.86 -0.73 0.95 -0.77 1.03 -0.72 0.90
Noise & Tactical 20 0.76 0.64 0.58 0.62 0.78 0.64 0.81 0.64
60 -0.25 0.67 -0.36 0.64 -0.23 0.68 -0.25 0.67
Noise & Tactical 20 0.99 2.2 0.69 2.73 1.09 2.28 1.13 2.08
& Strategic 60 0.00 2.22 0.11 2.17 0.10 1.96 0.23 2.15
(a) Noise Traders
(b) Noise & Tactical Traders
(c) Noise, Tactical & Strategic Traders
Figure 13: Reward densities of the different RL algorithms. LN corresponds to the logistic–normal algorithm that uses all features. LN-O corresponds to the algorithm without level and queue position features. LN-V corresponds to the algorithm without volume features. LN-D corresponds to the algorithm without the order flow imbalance and mid-price drift features. The upper panels correspond to an initial position of 20 lots. The lower panel corresponds to an initial position of 60 lots.